OpenStudy (anonymous):

Let f(x)=(6x^3-7)^3 and g(x)=6x^3-7 Given that f(x)=(h*g)(x), find h(x).

OpenStudy (anonymous):

So divide f(x) by g(x)

OpenStudy (anonymous):

and you will have h(x)

OpenStudy (anonymous):

\[f(x)/g(x)=h(x)\]

OpenStudy (haseeb96):

f(x) = (h*g)(x) f(x)=h(x)* g(x) f(x)/g(x)=h(x) now put the equation in it and solve for h(x).

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

\[(6x ^{3}-7)^{3}/(6x ^{3}-7)=( 6x ^{3}-7)^{2}\]

OpenStudy (anonymous):

and that is your h(x)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks to openstudy